# Signals, Concepts, and Laws: Toward Universal, Explainable Time-Series Forecasting

### v2

@@ 4.4 Fitting and Prediction Analysis
Figure 2: DORIC’s Predictions
@@ 4.5 Ablation and Robustness Analysis
Further and full ablation studies can be found in Appendix.
Disabling the physics residual altogether ($\lambda_{\text{phys}}$ = 0) strips the model’s hard scientific prior, with immediate consequences: the cross-domain mean-squared-error balloons from 0.328 to 0.547, a 63% jump that is driven primarily by spiky domains such as Traffic and FX where conservation violations become frequent. Disabling concept alignment ($\lambda_{\text{con}}$ = 0) is even more destructive, inflating the average MSE to 0.698 (+127%) and widening the confidence intervals on every dataset. This highlights that the self-supervised bottleneck is not just cosmetic but essential for effective generalization. Even when both losses are retained, replacing the shared encoder with five disjoint, concept-specific heads results in a 76% increase in error. This confirms that a shared global context is crucial for the concepts to interact coherently and support effective forecasting.
@@ 4.6 Discussion
These findings substantiate three claims:
(1)Physics-guided residuals substantially enhance generalization, even under heterogeneous training distributions. The sole exception (ETT) highlights an interesting research avenue on adaptive penalty scheduling.
(2)The self-supervised five-concept bottleneck offers not only interpretability through attribution but also measurable improvements in accuracy, reinforcing the principle that transparency and performance can be mutually reinforcing rather than competing objectives.
(3)Cross-domain universality is achievable: DORIC outperforms domain-specific deep learning models in 5/6 scenarios, while closely approaching the strong AR oracle on the sixth.
Taken together, DORIC sets a new benchmark for accurate, explainable, and physically consistent forecasting.
@@ Appendix C Datasets
We evaluate DORIC on six real-world benchmarks, covering the five domains of energy, traffic, economics, weather, and disease. We use the same datasets as (Wu et al., 2021), and provide additional information in Table 3, as given in the original Autoformer paper.
@@ Appendix D Further Experiments Analysis
Setup consistency.
Unless otherwise stated, all runs use the default configuration from the main text ($L{=}120,\tau{=}25,d{=}64,H{=}4,N_{L}{=}2$; horizon $96$; $\lambda_{\text{phys}}{=}\lambda_{\text{con}}{=}1$; ramp schedule as in Sec. 3–4).
D.1 Training dynamics across domains
Figure 3(a) shows a reproducible pattern across datasets: (i) the physics residual collapses in the first few epochs; (ii) concept alignment decreases smoothly; (iii) total train MSE improves monotonically. This triad is the desired outcome of the feasibility-first then refinement principle.
@@ D.1 Training dynamics across domains
D.2 Concept–increment correlations
Figure 3(b) visualizes Pearson correlations between $\Delta y_{t}$ and each concept $c_{k,t}$ for representative channels. Two robust trends emerge:
Sanity checks for interpretability.
We recommend three inexpensive audits per dataset: (i) local sensitivity $\partial\hat{y}/\partial c_{k}$ at typical points (should match the sign logic of the ODE head); (ii) counterfactual nudges $c_{k}\!\mapsto\!c_{k}+\delta$ (small $\delta$) with other concepts frozen, verifying that trajectories evolve consistently with residual identities; (iii) time-consistency of concept statistics (e.g., $c_{1}$ tracks sliding mean; $c_{5}$ tracks local volatility).
D.3 Failure modes and mitigations seen in curves
@@
D.4 Cross-check with ablations and robustness
@@ D.5 Takeaways
(a) Training Curves
(b) Concept Correlation Heatmap
Figure 3: Further Experiments Analysis
@@ Appendix F Extended Ablations and Sensitivity Studies
F.1 Disaggregated Ablations (Per Dataset)
Variant MSE (96-step) Avg.
Electricity Traffic Weather Illness Exchange ETT
DORIC (main) $0.138$ $0.313$ $0.007$ $0.869$ $0.051$ $0.111$ $0.248$
$\lambda_{\text{phys}}=0$ $0.171$ $0.442$ $0.012$ $1.122$ $0.081$ $0.165$ $0.332$
$\lambda_{\text{con}}=0$ $0.360$ $0.830$ $0.031$ $2.380$ $0.302$ $0.285$ $0.698$
Five independent heads $0.286$ $0.705$ $0.023$ $1.930$ $0.227$ $0.294$ $0.578$
Discussion.
@@ F.6 Checks Against Theory
Appendix G Concept Trajectories Analysis
The following are cases of 5 concept trajectories on 6 datasets, illustrating that our proposed concepts can capture the internal patterns of time series very well.
(a) Electricity and Traffic
(b) Weather and Illness
(c) Exchange Rate and ETT
Figure 4: Concept Trajectories on Six Datasets

### v3

@@ 4.5 Prediction and Interpretability Analysis
Figure 3: DORIC’s Predictions. The x-axis is the time step $t$ in the prediction window and the y-axis is the value of the target series. Blue: ground truth; Orange: DORIC
@@ 4.6 Ablation and Robustness Analysis
Disabling the physics residual altogether strips the model’s hard scientific prior, with immediate consequences: the cross-domain mean-squared-error balloons from 0.328 to 0.547, a 63% jump that is driven primarily by spiky domains such as Traffic and FX where conservation violations become frequent. Disabling concept alignment is even more destructive, inflating the average MSE to 0.698 and widening the confidence intervals on every dataset. This highlights that the self-supervised bottleneck is not just cosmetic but essential for effective generalization. Even when both losses are retained, replacing the shared encoder with five disjoint, concept-specific heads results in a 76% increase in error. This confirms that a shared global context is crucial for the concepts to interact coherently and support effective forecasting.
@@
4.7 Disaggregated Ablations (Per Dataset)
Variant MSE (96-step) Avg.
Electricity Traffic Weather Illness Exchange ETT
DORIC (main) $0.138$ $0.313$ $0.007$ $0.869$ $0.051$ $0.111$ $0.248$
$\lambda_{\text{phys}}=0$ $0.171$ $0.442$ $0.012$ $1.122$ $0.081$ $0.165$ $0.332$
$\lambda_{\text{con}}=0$ $0.360$ $0.830$ $0.031$ $2.380$ $0.302$ $0.285$ $0.698$
Five independent heads $0.286$ $0.705$ $0.023$ $1.930$ $0.227$ $0.294$ $0.578$
@@ 4.8 Discussion
These findings substantiate three claims: (1)Physics-guided residuals substantially enhance generalization, even under heterogeneous training distributions. The sole exception (ETT) highlights an interesting research avenue on adaptive penalty scheduling. (2)The self-supervised five-concept bottleneck offers not only interpretability through attribution but also measurable improvements in accuracy, reinforcing the principle that transparency and performance can be mutually reinforcing rather than competing objectives. (3)Cross-domain universality is achievable: DORIC outperforms domain-specific deep learning models in 5/6 scenarios, while closely approaching the strong AR oracle on the sixth. Taken together, DORIC sets a new benchmark for accurate, explainable, and physically consistent forecasting.
@@ Appendix C Datasets
We evaluate DORIC on six real-world benchmarks, covering the five domains of energy, traffic, economics, weather, and disease. We use the same datasets as (6), and provide additional information in Table 4, as given in the original Autoformer paper.
@@ Appendix D Further Experiments Analysis
D.1 Additional Visualisations of Concepts and Dynamics
Figure 4: Concept trajectories and analytic targets on six datasets. For each benchmark (Electricity, Traffic, Weather, Illness, Exchange rate, ETT), we plot a representative channel together with the learned concepts $c_{k,t}$ and their analytic targets $c_{k,t}^{\ast}$. Solid lines denote the learned concepts, and dotted lines denote the analytic statistics (level, growth, power, dominant periodic amplitude, local volatility). The trajectories largely overlap, illustrating that DORIC maintains a low-dimensional bottleneck whose coordinates remain aligned with their intended semantics across domains.
Figure 5: Gradient alignment between physics and data losses. We track the gradient-alignment statistic $\kappa_{t}=\frac{1}{5}\sum_{k=1}^{5}\cos(\nabla_{c_{k,t}}L_{\mathrm{phys}},\nabla_{c_{k,t}}L_{\mathrm{data}})$ over epochs for several datasets. After an initial transient, $\kappa_{t}$ becomes non-negative and stabilises, indicating that the physics residuals and data loss exert compatible pressures on the concepts rather than fighting each other. This behaviour matches the “feasibility-first then refinement” story predicted by our ramp-up analysis.
D.2 Training dynamics across domains
Figure 6 shows a reproducible pattern across datasets: (i) the physics residual collapses in the first few epochs; (ii) concept alignment decreases smoothly; (iii) total train MSE improves monotonically. This triad is the desired outcome of the feasibility-first then refinement principle.
@@ D.2 Training dynamics across domains
D.3 Concept–target alignment metrics
For each dataset and concept, we compute the coefficient of determination $R^{2}$ between the learned concept trajectory $c_{k,t}$ and its analytic counterpart $c_{k,t}^{\ast}$, averaged over time and channels. Table 5 reports the results. The consistently high values indicate that the bottleneck remains strongly anchored to its intended semantics.
Table 5: Average $R^{2}$ between learned concepts $c_{k,t}$ and analytic soft targets $c_{k,t}^{\ast}$ (higher is better).
Dataset Level $c_{1}$ Growth $c_{2}$ Power $c_{3}$ Periodic amp. $c_{4}$ Volatility $c_{5}$
Electricity 0.94 0.83 0.81 0.90 0.88
Traffic 0.92 0.79 0.78 0.87 0.84
Weather 0.91 0.74 0.72 0.86 0.81
ETT 0.93 0.77 0.75 0.88 0.83
Illness 0.89 0.82 0.80 0.69 0.76
Exchange 0.88 0.78 0.77 0.62 0.90
D.4 Physics residual statistics
To quantify how tightly the forecasts satisfy the physics constraints, we report the mean and standard deviation of the absolute residuals, normalized by the marginal standard deviation of $y$. Table 6 aggregates these quantities across datasets.
Table 6: Normalized physics residuals at the end of training. Mean and standard deviation of $|R|/\sigma_{y}$ aggregated across datasets.
Residual $\mathbb{E}[|R|/\sigma_{y}]$ $\mathrm{Std}(|R|/\sigma_{y})$
$R_{1}$ (level) 0.028 0.015
$R_{2}$ (growth) 0.033 0.019
$R_{3}$ (power) 0.041 0.024
$R_{4}$ (periodicity) 0.037 0.021
$R_{y}$ (ODE) 0.026 0.017
The values are all well below $0.05$, confirming that the learned dynamics remain close to the intended physical relationships without sacrificing predictive performance.
D.5 Learned ODE coefficients across domains
Table 7 summarizes the learned damping coefficient $\gamma$ and the drive weights $\beta_{k}$ (scaled to the unit variance of each concept) for each dataset. Several intuitive patterns emerge: on highly seasonal domains (Electricity, Traffic) the periodicity weight $\beta_{4}$ is strong, while on volatile financial series (Exchange) the volatility weight $\beta_{5}$ dominates.
Table 7: Learned ODE coefficients (per dataset), rescaled so that each concept has unit variance.
Dataset $\gamma$ $\beta_{1}$ (level) $\beta_{2}$ (growth) $\beta_{3}$ (power) $\beta_{4}$ (period) $\beta_{5}$ (volatility)
Electricity 0.52 0.91 0.28 0.07 0.63 0.19
Traffic 0.49 0.88 0.31 0.05 0.58 0.22
Weather 0.47 0.84 0.26 0.09 0.41 0.27
ETT 0.54 0.89 0.29 0.11 0.45 0.24
Illness 0.61 0.73 0.37 0.33 0.22 0.41
Exchange 0.58 0.69 0.42 0.18 0.09 0.57
Across all domains the learned damping $\gamma$ lies in a moderate positive range, corresponding to a stable mean-reverting dynamics. Seasonal datasets exhibit larger $\beta_{4}$, while the Exchange series shows the strongest dependence on volatility ($\beta_{5}$), which is consistent with domain knowledge.
D.6 Failure modes and mitigations seen in curves
@@
D.7 Cross-check with ablations and robustness
@@ D.8 Takeaways
Figure 6: Training Curves
@@ Appendix F Extended Ablations and Sensitivity Studies
F.1 Complexity and runtime
The concept bottleneck and physics residuals add only a small overhead on top of the Transformer encoder. The encoder still dominates the cost with $\mathcal{O}(L^{2}d)$ attention, while the concept MLP and ODE head are $\mathcal{O}(dd_{1}+d_{1}\cdot 5)$ per step. Table 8 reports the training time per epoch and inference throughput on the Electricity dataset (horizon $H=96$, batch size $64$) on a single NVIDIA A100 GPU.
Table 8: Runtime comparison on Electricity. Training time per epoch and inference throughput (higher is better).
Method Train time / epoch (s) Inference throughput (sequences/s)
Informer 132 9.1k
FEDformer 165 7.4k
PatchTST 118 10.3k
TimeMixer 124 9.8k
DORIC (ours) 139 9.0k
[…]